OpenStudy (kingj):

In a ping pong tournament, 30% of the players are eliminated each day. If the tournament starts with 2,000 players, to the nearest integer, how many players will be present at the start of the 5th day of play? 259 480 819 1,229

OpenStudy (anonymous):

i got 480. .

OpenStudy (anonymous):

initial number= 2000 first day - 2000(.3) =600 will be eliminated therefor; 2000-600 = 1400, can proceed to the next round. second day- 1400(.3) =420 will be eliminated therefor; 1400-420= 980, can proceed to the next round third day- 980(.3) =294 will be eliminated therefor; 980-296= 686,can proceed to the next round fourth day- 686(.3)=205.8 equivalent to 206 will be eliminated therefor; 686-206=480 can proceed to the next round the final answer is 480, will be present in the start of the 5th day. .